# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@
$\hat{\theta}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}=\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right\}.$ (2.3)
We call $\hat{\theta}_{\textsc{MAP}}$ maximum a posteriori estimator, or shortly MAP estimator.
@@
    Up to now, we do not have formulated any specific conditions on $\tilde{\Theta}$, except of $\tilde{\Theta}\subseteq\mathbb{L}^{2}(\mathcal{O},W)$. In Proposition 2.5, a restriction of $\tilde{\Theta}$ is neccessary in order to provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer of the optimization problem Eq. 2.3. However, the results of the statistical analysis hold true for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$. From the application perspective, the definition of $\hat{\theta}_{\textsc{MAP}}$ seems unfeasible as the optimization problem is formulated over a possibly infinite-dimensional space. The theory shown in Nickl [2023] and Nickl and Wang [2024] provide general conditions and algorithms, under which $\hat{\theta}_{\textsc{MAP}}$ can be computed in polynomial time, if $\tilde{\Theta}$ is finite dimensional.
@@
    The MAP estimator $\hat{\theta}_{\textsc{MAP}}$ introduced in Eq. 2.3 has two different interpretations: On the one hand, $\hat{\theta}_{\textsc{MAP}}$ is strongly related to the classical theory of Tikhonov-regularized estimators - we refer to Engl et al. [1996] for a detailed analysis of the regularization of ill-posed deterministic inverse problems. Indeed, $\hat{\theta}_{\textsc{MAP}}$ minimizes the empirical squared distance. To exlude unnecessary complex solutions of the problem, the regularity is penalized in terms of the $H^{\alpha}(\mathcal{O},W)$-norm. On the other hand, $\hat{\theta}_{\textsc{MAP}}$ corresponds to the framework of Bayesian inverse problem techniques. Under Gaussian process priors on the parameter space $\tilde{\Theta}$, the corresponding likelihood of the a posteriori distribution is then maximised by $\hat{\theta}_{\textsc{MAP}}$. We refer here to the comments in Nickl et al. [2020] and Bohr and Nickl [2023], and further to Dashti et al. [2013], Helin and Burger [2015] for more details.
   A convergence rate result
The main results of this work - Proposition 2.5, Theorem 2.6 and Corollary 2.7 - are presented in the following. Firtsly, we provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a solution to the optimization problem Eq. 2.3 over a suitable subset $\Theta\subseteq\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$. Moreover, we are showing a concentration result controlling the prediction error and the regularity of $\hat{\theta}_{\textsc{MAP}}$ simultanously under $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability. Therefore, we define for every $\theta_{1}\in\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$, $\theta_{2}\in\tilde{\Theta}$, and $\mathfrak{r}\in\mathbb{R}_{\geq 0}$ the functional
$\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{1},\theta_{2}):=\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
@@ 1. C1)
    $\mathbb{L}^{2}$-Lipschitz continuity: Let $\alpha,\gamma_{1},\kappa\in\mathbb{R}_{\geq 0}$. Set $\mathscr{H}=H^{\alpha}_{c}(\mathcal{O},W),\text{ if $\kappa\geq\frac{1}{2}$ }$ and $\mathscr{H}=H^{\alpha}(\mathcal{O},W),\text{ if $\kappa<\frac{1}{2}$}$. The map $\mathscr{G}$ is $(\alpha,\gamma_{1},\kappa)$-regular in $\mathbb{L}^{2}$-norm, if there exists a constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in\tilde{\Theta}\cap\mathscr{H}$
    $\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq C_{\operatorname{Lip},2}\cdot\left(1+\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\lor\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
@@ 2. C2)
    $\sup_{\theta\in\tilde{\Theta},z\in\mathcal{Z}}\left\{\left|\mathscr{G}(\theta)(z)\right|_{V}\right\}\leq C_{\mathscr{G},\operatorname{B}}.$
@@ 3. C3)
    $\mathbb{L}^{\infty}$-Lipschitz continuity: Let $\alpha,\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$, such that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. The map $\mathscr{G}$ is $(\alpha,\gamma_{2},\kappa)$-regular in $\mathbb{L}^{\infty}$-norm, if there exists a constant $C_{\operatorname{Lip},\infty}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O},W)$
@@
    $\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\infty}\leq C_{\operatorname{Lip},\infty}\cdot\left(1+\left\|\theta_{1}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\left\|\theta_{2}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Remark § 2.4 ($\mathbb{L}^{\infty}$-Lipschitz continuity).
The effect that stronger assumptions on the statistical model are needed to prove consistency for the maximum likelihood estimator than to show corresponding concentration results in the Bayesian setting (see Le Cam [1986]) is also apparent in the nonlinear random regression setting considered here: While in Nickl [2023] it suffices to assume condition C1) and C2) (as special case with $\gamma_{1}=0$) to obtain posterior convergence results, in Bohr and Nickl [2023] consistency of the MAP estimator necessitates a $\mathbb{L}^{\infty}$-stability estimate of the forward map $\mathscr{G}$, i.e.,
$\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\infty}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\infty},$
uniformly over the whole parameter space. C3) in 2.3 generalizes this condition by replacing the uniform norm by the stronger $\left\|\cdot\right\|_{C^{\eta_{2}}}$-norm. Moreover, we now allow a polynomial increase of order $\gamma_{2}$. While the conditions formulated in Bohr and Nickl [2023] (which correspond to the special case $\eta_{1}=\eta_{2}=\gamma_{2}=0$ in the present setting!) are insufficient for the Darcy problem, we see in Section 3 and Appendix A that the associated solution map fulfills C3) with $\eta_{1}=\eta_{2}=1$ and $\gamma_{2}=4$. Note that the condition $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$ in C3) ensures that all norms are well-defined as in this situation the embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{i}}(\mathcal{O},W)$, $i\in\left\llbracket 2\right\rrbracket$, holds true.
@@
The first main results addresses the open question of existence of $\hat{\theta}_{\textsc{MAP}}$ as the maximizer of the penalized Tikhonov-regularized functional Eq. 2.2. For the white noise model, the authors in Nickl et al. [2020] provide a proof using the white noise formulation of the regression model. We adapt the corresponding proof to the random design regression model, also exploiting the Direct Method of the Calculus of Variations, making reference to Dacorogna [2008]. The proof can be found in Appendix A.
@@
for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, some reference parameter $\tilde{\theta}\in\Theta$ (with $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$) and some $c_{1}\in\mathbb{R}_{\geq 1}$. To shorten notation, we define the family of symbols $\texttt{P}:=\left\{\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\sigma\right\}$ and write $c(\texttt{P})$ for a constant to emphasize the dependency.
@@
Suppose that $\mathscr{G}:\tilde{\Theta}\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfies C1)-C3) in 2.3 with some $\alpha>\max\left\{\eta_{1}\lor\eta_{2}+d\lor\frac{d}{2}(1+\gamma_{2}),\frac{d}{2}\gamma_{1}-\kappa\right\}$. Let $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ from Eq. 2.1 with some fixed $\theta_{\mathfrak{o}}\in\tilde{\Theta}$. Then the following hold.
@@
    $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{2}\cdot\exp(-\bar{c}\cdot NR^{2})$ (2.4)
@@
    $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}\left[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\right]\leq c_{3}\cdot\inf_{\tilde{\theta}\in\Theta}\left\{\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}(1+\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O})}^{2})+\frac{1}{N}\right\}.$ (2.5)
@@
Under the assumptions of Theorem 2.6, assume $\theta_{\mathfrak{o}}\in\Theta$ with $\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}\in\mathbb{R}_{>0}$. Let $\mathfrak{r}\in\mathbb{R}_{>0}$ be given by
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{4}\mathfrak{r}^{2}_{N}\right\}\right)\leq c_{5}\exp(-\bar{c}N\mathfrak{r}_{N}^{2})$
@@
$\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}\left[\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta)\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right]\right\}\leq\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta)]\right\}\leq c_{6}\cdot\mathfrak{r}_{N}^{2}.$
@@
    Note, that Theorem 2.6 does not require $\theta_{\mathfrak{o}}\in\Theta$ as long as it can be approximated sufficiently by some $\tilde{\theta}\in\Theta$. Therefore, the assumption $\theta_{\mathfrak{o}}\in\Theta$ in Corollary 2.7 can also be weakened to the effect that $\mathscr{G}(\tilde{\theta})$ approximates $\mathscr{G}(\theta_{\mathfrak{o}})$ in some sense (see equation 2.23 in Bohr and Nickl [2023] for instance).
@@
    Evidently, Corollary 2.7 provides the consistency of the MAP estimator $\hat{\theta}_{\textsc{MAP}}$. The rate of convergence for the prediction loss is known to be optimal for $\kappa$-smoothing linear forward maps $\mathscr{G}$.
Without any further assumptions, there is no hope to obtain an analogous upper bound for the estimation risk as well. However, the next corollary demonstrates stability estimates under a inverse continuity modulus of $\mathscr{G}^{-1}$.

### v3

@@ 2.3 Maximum a posteriori estimator
$\hat{\theta}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}.$ (9)
@@
    Up to now, we do not have formulated any specific conditions on $\tilde{\Theta}$, except of $\tilde{\Theta}\subseteq\mathbb{L}^{2}(\mathcal{O},W)$. In proposition 2.5, we need to restrict ourselves to suitable subsets $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$ (with $\mathscr{H}$ being either $H^{\alpha}(\mathcal{O},W)$ or $H^{\alpha}_{c}(\mathcal{O},W)$) in order to provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer of the optimization problem eq. 9 over $\Theta$. However, the results of the statistical analysis hold true for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$. From the application perspective, the definition of $\hat{\theta}_{\textsc{MAP}}$ seems unfeasible as the optimization problem is formulated over a possibly infinite-dimensional space. The theory shown in Nickl (2023) and Nickl and Wang (2024) provide general conditions and algorithms, under which $\hat{\theta}_{\textsc{MAP}}$ can be computed in polynomial time, if $\Theta$ is finite dimensional.
@@
    The estimator $\hat{\theta}_{\textsc{MAP}}$ introduced in eq. 9 admits two different but related interpretations. On the one hand, it is closely connected to the classical theory of Tikhonov-regularized estimators, as studied in Engl et al. (1996), where the regularization of ill-posed deterministic inverse problems is achieved by penalizing solutions with respect to a suitable Sobolev norm. Specifically, $\hat{\theta}_{\textsc{MAP}}$ minimizes the empirical squared loss subject to a regularization term given by the $H^{\alpha}(\mathcal{O},W)$-norm, thereby enforcing smoothness to exclude overly complex solutions. On the other hand, $\hat{\theta}_{\textsc{MAP}}$ corresponds to the maximum a posteriori (MAP) estimator in a Bayesian inverse problem framework: Let $\Pi$ be a Gaussian process prior on the parameter $\theta$ with reproducing kernel Hilbert space (RKHS) $\mathscr{H}$ and corresponding prior density $\pi(\theta)\propto\exp(-\frac{\overline{\mathfrak{r}}^{2}}{2}\|\theta\|_{H^{\alpha}(\mathcal{O},W)}^{2})$ with $\overline{\mathfrak{r}}>0$. Under the Gaussian likelihood function of the model eq. 7, any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\mathscr{H}$ with $\mathfrak{r}=\frac{1}{\sqrt{N}}\overline{\mathfrak{r}}$ has a formal interpretation as maximizer of the log-likelihood $\log\pi(\theta|D_{N})$ of the resulting posterior distributions $\Pi(\cdot|D_{N})$. This interpretation has been formalized, for instance, in Remark 4 of Nickl et al. (2020). We refer also to Bohr and Nickl (2023); Dashti et al. (2013), and Helin and Burger (2015) for additional background on MAP estimation in Gaussian Bayesian inverse problems. Throughout this work, we refer to $\hat{\theta}_{\textsc{MAP}}$ as the maximum a posteriori (MAP) estimator.
2.4 A convergence rate result
The main results of this work - proposition 2.5, theorem 2.6 and corollary 2.7 - are presented in the following. Firstly, we provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a solution to the optimization problem eq. 9 over a suitable subset $\Theta\subseteq\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$. Moreover, we are showing a concentration result controlling the prediction error and the regularity of $\hat{\theta}_{\textsc{MAP}}$ simultanously under $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability. Therefore, we define for every $\theta_{1}\in\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$, $\theta_{2}\in\tilde{\Theta}$, and $\mathfrak{r}\in\mathbb{R}_{\geq 0}$ the functional
$\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{1},\theta_{2}):=\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}\|\theta_{1}\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
@@ 1. C1)
[…]
    $\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq C_{\operatorname{Lip},2}\left(1+\|\theta_{1}\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\hskip-4.2679pt\lor\|\theta_{2}\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\right)\|\theta_{1}-\theta_{2}\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
@@ 2. C2)
    $\sup_{\theta\in\tilde{\Theta},z\in\mathcal{Z}}\left\{|\mathscr{G}(\theta)(z)|_{V}\right\}\leq C_{\mathscr{G},\operatorname{B}}.$
@@ 3. C3)
    $\mathbb{L}^{\infty}$-Lipschitz continuity: Let $\alpha,\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$, such that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. The map $\mathscr{G}$ is $(\alpha,\gamma_{2},\eta_{1},\eta_{2})$-regular in $\mathbb{L}^{\infty}$-norm, if there exists a constant $C_{\operatorname{Lip},\infty}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in\tilde{\Theta}\cap\mathscr{H}$
@@
    $\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\infty}\leq C_{\operatorname{Lip},\infty}\left(1+\|\theta_{1}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\|\theta_{2}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\|\theta_{1}-\theta_{2}\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Remark § 2.4 ($\mathbb{L}^{2}$- and $\mathbb{L}^{\infty}$-Lipschitz continuity).
The assumptions C1)–C3) imposed on the forward map $\mathscr{G}$ are abstract but serve as minimal requirements for establishing statistical guarantees for the MAP estimator $\hat{\theta}_{\textsc{MAP}}$. Condition C1) quantifies the degree of ill-posedness through the smoothing parameter $\kappa$, which also directly influences the convergence rates obtained in theorem 2.6. In contrast, C3) is of a more technical nature: it ensures that entropy numbers of images $\{\mathscr{G}(\theta):\theta\in H^{\alpha}(\mathcal{O},W,r)\}$, $r\in\mathbb{R}_{>0}$, of Sobolev balls under $\mathscr{G}$ are sufficiently controlled and is particularly tailored to the PDE-constrained setting considered here, such as the Darcy flow. For this purpose, we weaken the $\mathbb{L}^{\infty}$-stability condition
$\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\|_{\infty}\lesssim\|\theta-\theta^{\prime}\|_{\infty},$ (10)
as used in Bohr and Nickl (2023), by replacing the uniform norm on the right with the stronger $C^{\eta_{2}}$-norm and additionally allowing polynomial growth of order $\gamma_{2}$. The parameters $\eta_{1}$ and $\eta_{2}$ do not appear in the convergence rates obtained in theorem 2.6 explicitly, but the condition $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$ ensures that the relevant embeddings $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{i}}(\mathcal{O},W)$, $i\in\{1,2\}$, hold, so that all norms are well-defined. It is also not surprising that stronger assumptions are needed to establish consistency of the MAP estimator than to prove posterior contraction in the Bayesian setting, such as in Giordano and Nickl (2020) for the Darcy problem. This phenomenon is well known in classical statistics (see, e.g., Le Cam (1986)) and becomes evident in the general nonlinear random regression framework as well: while Nickl (2023) derives posterior contraction under the comparatively mild conditions C1) and C2) (with $\gamma_{1}=0$), the analysis in Bohr and Nickl (2023) requires the uniform stability condition eq. 10 to prove consistency of the MAP estimator. In our setting, the condition eq. 10 correspond to the special case $\eta_{1}=\eta_{2}=\gamma_{2}=0$, which is insufficient for the Darcy problem. As we demonstrate in section 3 and section 4, the associated solution map fulfills C3) with $\eta_{1}=\eta_{2}=1$ and $\gamma_{2}=4$.
@@
The first main results address the open question of existence of $\hat{\theta}_{\textsc{MAP}}$ as the maximizer of the penalized Tikhonov-regularized functional eq. 8. For the white noise model, the authors in Nickl et al. (2020) provide a proof using the white noise formulation of the regression model. We adapt the corresponding proof to the random design regression model, also exploiting the Direct Method of the Calculus of Variations, making reference to Dacorogna (2008). The proof can be found in section 4.
@@
for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, some reference parameter $\tilde{\theta}\in\Theta$ (with $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$) and some $c_{1}\in\mathbb{R}_{\geq 1}$. To shorten notation, we define the family of symbols
$\texttt{P}:=\left\{\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\sigma\right\}$
and write $c(\texttt{P})$ for a constant to emphasize the dependency.
@@
Suppose that $\mathscr{G}:\tilde{\Theta}\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfies C1)-C3) in 2.3 with some $\alpha>\max\left\{\eta_{1}\lor\eta_{2}+d\lor\frac{d}{2}(1+\gamma_{2}),\frac{d}{2}\gamma_{1}-\kappa\right\}$. Let $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ from eq. 7 with some fixed $\theta_{\mathfrak{o}}\in\tilde{\Theta}$. Then the following hold.
@@
    $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{2}\cdot\exp(-\bar{c}\cdot NR^{2})$ (11)
@@
    $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}\left[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\right]\leq c_{3}\cdot\inf_{\tilde{\theta}\in\Theta}\left\{\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}(1+\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O})}^{2})+\frac{1}{N}\right\}.$
@@
Under the assumptions of theorem 2.6, let $\theta_{\mathfrak{o}}\in\Theta$ with $\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}\in\mathbb{R}_{>0}$. Let $\mathfrak{r}\in\mathbb{R}_{>0}$ be given by
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{4}\mathfrak{r}^{2}_{N}\right)\leq c_{5}\exp(-\bar{c}\cdot N\mathfrak{r}_{N}^{2})$
@@
$\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}\left[\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta)\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right]\right\}\leq c_{6}\cdot N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}.$
@@
[…]
@@
    Evidently, corollary 2.7 provides the consistency of the MAP estimator $\hat{\theta}_{\textsc{MAP}}$. The rate of convergence for the prediction loss fits the interpretation of classical inverse problems: The smoothness parameter $\kappa$ of the forward map $\mathscr{G}$, quantifying the ill-posedness of the problem, influences directly the rate of convergence. It is known that this rate of convergence is minimax-optimal for $\kappa$-smoothing linear forward maps $\mathscr{G}$ in various settings, see, e.g., Cavalier (2008). In contrast, $\eta_{1}$ and $\eta_{2}$ from condition C3) do not appear explicitely in the rate, but ensure by the condition $\alpha>\eta_{1}\lor\eta_{2}+d\lor\frac{d}{2}(1+\gamma_{2})$ that the model is sufficiently regular.
Without any further assumptions, there is no hope to obtain an analogous upper bound for the estimation risk as well. However, the next corollary demonstrates stability estimates under an inverse continuity modulus of $\mathscr{G}^{-1}$.